OpenStudy (anonymous):

(MEDALS!!!!!!!!) Please answer. Group the terms with variables on one side of the equal sign, and simplify. -7s + 12 = -9s

OpenStudy (anonymous):

Add 7s to both sides first

OpenStudy (anonymous):

Add 7s to both sides 12 = -2s Divide by -2 on both sides S = -6 Correct?

OpenStudy (pinksapphire):

-7s + 12 = -9s +7s +7 12=-2s 12/-2=-2x/-2 -6=s

OpenStudy (anonymous):

That's right, good job